The length of a rectangular swimming pool is twice its width. The pool is surrounded by a walk that is 2 feet wide. The area of the region consisting of the pool and the walk is 1056 square feet.
(a) Use the method of completing the square to determine the dimensions of the swimming pool.
The length of the pool is ____ feet, and the width of the pool is ____feet.
(b) If the material for the walk costs $8 per square foot, how much would the material cost for the entire walk?
The material would cost $____ .

Answer by jorel555(1290)

First, l=2w. The dimensions of the area are w+4 and 2w+4. So:
a.(w+4)(2w+4)=1056
2w^2+12w+16=1056
w^2+6w-520=0
(w+26)(w-20)=0
w=20,-26
The width of our pool is 20 ft., and it's length is 40 ft..
b.The total area (1056) less the area of the pool (800) equals 256 sq. ft.. At $8 per square foot, that amounts to $2048 total dollars for the material!!!!!!!!!!